# Effective algorithms for tensor train decomposition via the UTV framework

### v1

@@ Abstract
The tensor train (TT) decomposition is used to compress large tensors into a more compact form by exploiting their inherent data structures. A fundamental approach for constructing the TT format is the TT-SVD, which extracts the TT-cores by the singular value decompositions (SVDs) sequentially. But in practical applications, it is often not necessary to compute full SVDs. In this article, we therefore propose a new method called the TT-UTV. It utilizes the virtues of rank-revealing UTV decomposition to compute the TT format for a large-scale tensor, hence requires less computational cost. We analyze the error bounds on the accuracy of these algorithms both in the URV and ULV cases, then recommend different sweep patterns for these two cases. We perform numerical experiments on some applications, including magnetic resonance imaging (MRI) data completion, to illustrate their good performance in practice.
Index Terms:
@@
Tensor train decomposition, UTV decomposition, tensor completion.
I Introduction
@@
Real-world high-dimensional data are usually represented as tensors, such as genetic analysis, video sequences, hypergraphs, hyperspectral images, and social network interactions [14, 23, 5, 41, 21]. Tensors provide a natural way to model multi-dimensional data and multi-modal nonlinear interactions. Due to the “curse of dimensionality” [15], direct usage of these higher-order tensors requires a lot of resources as the time and space complexity grow exponentially with the order of the tensor. Fortunately, tensors in the real world often have low-rank structure and the tensor-train decomposition [29, 28], also well-known as matrix product state representation [30], can exploit the inherent low-rank structure to mitigate the curse of dimensionality in many practical scenarios.
The tensor train (TT) decomposition has found widespread applications in areas such as machine learning [44, 27, 34], quantum many-body physics [32], signal processing [33], data completion [35, 2] and numerical partial differential equations [8, 31]. For a $d$th-order tensor, the TT-SVD algorithm [29] sequentially utilizes $d-1$ singular value decompositions to extract core tensors one by one, which becomes a foundational approach to decompose a higher-order tensor into the TT format.
@@
In many circumstances, the SVD is computationally demanding and difficult to update, which gives birth to the UTV decomposition [36]. It provides approximately the same numerical rank and subspace information compared to SVD, while maintaining significantly lower computational cost [36, 10, 9, 7]. Moreover, UTV decomposition can be implemented to achieve higher performance on modern hardware architectures by leveraging block matrix operations, which are well suited for parallel processing [26]. Due to its virtues, UTV decomposition becomes a common alternative for SVD in many applications including signal processing and machine learning [43, 45]. Recently, the UTV decomposition is also extended to the decompositions for higher-order tensors. For example, a truncated multilinear UTV decomposition [39] is introduced for tensors from the truncated multilinear SVD [6, 40], the t-UTV decomposition [4] is proposed for third-order tensors as a substitute to the t-SVD based on the t-product [17].
@@
In this article, we investigate whether the UTV decomposition can be leveraged to construct the TT decomposition, as economic alternatives to the TT-SVD. The goal is to save the computational cost by utilizing the advantages of UTV decomposition over SVD. Hence, a variety of algorithms for UTV decomposition can be employed for TT decomposition. This new approach is called TT-UTV. There exist two types of TT-UTV algorithms, corresponding to the URV and ULV cases, respectively. We provide an in-depth error analysis for these algorithms and demonstrate their effectiveness through various numerical experiments including image compression and MRI data completion.
@@
This paper is organized as follows. We report the involved notions in Section 2. In Section 3, we introduce the TT-ULV algorithm in left-to-right sweep and TT-URV algorithm in right-to-left sweep, respectively, and derive rigorous error bounds for them, which explain how the UTV decompositions are utilized to achieve accurate and efficient tensor-train decomposition. In Section 4, we conduct extensive performance evaluations of the TT-UTV algorithm with TT-SVD in various applications.
II Related works
@@
Notations. A $d$th-order tensor is a multidimensional array with $d$ free indices, matrices are second-order tensors and vectors are first-order tensors. We mainly inherit the usage of notations from the comprehensive works of Kolda et al. [20]. The calligraphic capital letters such as $\mathcal{A}$, boldface capital letters such as $\mathbf{A}$, and boldface lowercase letters such as $\mathbf{a}$ are used to denote the tensors, matrices, and vectors, respectively. $\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$ denotes the set of all $d$th-order tensors of dimensions $(I_{1},\dots,I_{d})$ over the real number field $\mathbb{R}$, where $I_{k}$ is the dimension of the $k$th mode of the tensor. The entries of $\mathcal{A}\in\mathbb{C}^{I_{1}\times\cdots\times I_{d}}$ are accessed by $\mathcal{A}_{i_{1}i_{2}\cdots i_{d}}$ or $\mathcal{A}(i_{1},i_{2},\cdots,i_{d})$ for $1\leq i_{k}\leq I_{k}$. For a tensor $\mathcal{A}$, we denote by $\mathbf{A}_{k}$ its $k$th unfolding matrix.
II-A Tensor basics
@@
Tensor unfolding, the process of rearranging the elements of a tensor into a matrix or vector, is a frequently used operation in tensor analysis and numerical computations [1, 20]. The reverse lexicographic ordering (as the ‘reshape’ command in MATLAB) is one of the most commonly used approaches for integrating multiple indices into a single long index. Concretely, for the tensor dimensions $\mathbf{I}=\left(I_{1},\dots,I_{d}\right)$ and a corresponding index vector $\mathbf{i}=\left(i_{1},\dots,i_{d}\right)$, define the index mapping
$\mathrm{ivec}(\mathbf{i},\mathbf{I}):=i_{1}+\sum_{k=2}^{d}(i_{k}-1)\prod_{j=1}^{k-1}I_{j}.$
Definition II.1 (The $k$th unfolding matrix [29]).
@@
For a $d$th order tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$, the matrix $\mathbf{A}_{k}\in\mathbb{R}^{(I_{1}\cdots I_{k})\times(I_{k+1}\cdots I_{d})}$ is called the $k$th unfolding matrix of $\mathcal{A}$ such that
@@
where $i=\mathrm{ivec}\left((i_{1},\dots,i_{k}),(I_{1},\dots,I_{k})\right)$ and $j=\mathrm{ivec}\left((i_{k+1},\dots,i_{d}),(I_{k+1},\dots,I_{d})\right)$.
The tensor train (TT) decomposition represents a higher-order tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$ as the contractions of small factor tensors
where $r_{0}=r_{d}=1$, these third-order tensors $\mathcal{G}^{(k)}\in\mathbb{C}^{r_{k-1}\times I_{k}\times r_{k}}$ for $k=1,2,\ldots,d$ are called TT-cores, and $(r_{1},r_{2},\ldots,r_{d-1})$ are TT-ranks. The TT-SVD algorithm computes the tensor format (1) with left-orthogonal cores by a single left-to-right sweep or right-orthogonal cores by a right-to-left sweep; more details can be found in [29].
Definition II.2 (Orthogonal cores[29]).
The cores $\{\mathcal{G}^{(k)}\}_{k=1}^{d}$ in (1) are called left-orthogonal if
@@
where $\mathbf{G}^{(k)}_{2}\in\mathbb{R}^{r_{k-1}I_{k}\times r_{k}}$ are the second unfolding matrices of the TT-cores $\mathcal{G}^{(k)}$. They are right-orthogonal if
$\mathbf{G}^{(k)}_{1}\mathbf{G}^{(k)\top}_{1}=\mathbf{I}_{r_{k-1}}$
where $\mathbf{G}^{(k)}_{1}\in\mathbb{R}^{r_{k-1}\times I_{k}r_{k}}$ are the first unfolding matrices of the TT-cores $\mathcal{G}^{(k)}$.
Definition II.3 (Mode product [20]).
The mode-$k$ product of $\mathcal{X}\in\mathbb{C}^{I_{1}\times\cdots\times I_{d}}$ with a matrix $\mathbf{A}\in\mathbb{C}^{J\times I_{k}}$ is an $I_{1}\times\cdots\times I_{k-1}\times J\times I_{k+1}\times\cdots\times I_{d}$ tensor, given by
@@
Apart from the tensor-train decomposition, the Tucker format [38] is another compact form for larger-scale tensors, which represents a $d$th-order tensor via the mode products of a smaller-scaled core tensor and $d$ factor matrices:
@@
For $\mathbf{A}\in\mathbb{C}^{m\times n}$ and $\mathbf{B}\in\mathbb{C}^{p\times q}$, the Kronecker product $\mathbf{A}\otimes\mathbf{B}$ is the $mp\times nq$ block matrix
@@ II-B UTV decomposition
where $\mathbf{R}_{11}\in\mathbb{R}^{r\times r}$, $\mathbf{R}_{22}\in\mathbb{R}^{(m-r)\times(n-r)}$ are upper triangular, $\mathbf{O}$ denotes the zero matrix whose elements are all 0, and $\mathbf{R}_{12}\in\mathbb{R}^{r\times(n-r)}$ with small entries. Moreover, suppose the numerical rank of $\mathbf{A}$ is $r$, then the UTV decomposition is said to be rank-revealing if
$\sigma_{\rm min}(\mathbf{R}_{11})=O(\sigma_{r})\quad\mathrm{and}\quad\left\|\begin{bmatrix}\mathbf{R}_{12}^{\top}&\mathbf{R}_{22}^{\top}\end{bmatrix}\right\|_{2}=O(\sigma_{r+1}),$ (4)
where $\sigma_{r}$ is the $r$th largest singular value of $\mathbf{A}$.
$\mathbf{A}=\mathbf{U}\begin{bmatrix}\mathbf{L}_{11}&\mathbf{O}_{r\times(n-r)}\\ \mathbf{L}_{21}&\mathbf{L}_{22}\end{bmatrix}\mathbf{V}^{\top},$ (5)
here $\mathbf{L}_{11}$ and $\mathbf{L}_{22}$ are lower triangular, $\mathbf{L}_{21}\in\mathbb{R}^{(m-r)\times r}$, and the ULV decomposition is said to be rank-revealing if
The smaller the norm of the off-diagonal block, the better the truncated low-rank approximations in (4) and (6). In fact, the UTV decomposition of a matrix is not unique, and various matrix decompositions proposed in the literature can be regarded as UTV-type, such as the rank-revealing QR decomposition [3], the pivoted QLP decomposition [37], and the SVD is also a special case of UTV decomposition when the middle matrix is diagonal. There are a lot of efficient algorithms for UTV decomposition. Stewart [36] introduced the algorithms for UTV decomposition that are well-suited for high-rank matrices. Fierro $et\ al.$ [12, 10] proposed the low-rank UTV algorithms and built up a Matlab package named ‘UTV Tools’ for efficient rank-revealing decompositions. In recent years, more fast randomized algorithms are proposed for UTV decomposition which are flexible for parallel processing in modern architectures [26, 22, 16]. In many practical applications, they can achieve comparable accuracy at a lower computational cost than the classical SVD.
III TT-UTV algorithms
In this section, we propose the TT-UTV algorithms for computing the approximate TT-format by sequentially separating the cores by the UTV decompositions, and demonstrate their feasibility through theoretical analysis.
III-A TT-ULV case

### v2

# Effective algorithms for tensor-train decomposition via the UTV framework

@@ Abstract
The tensor-train (TT) decomposition is widely used to compress large tensors into a more compact form by exploiting their inherent data structures. A fundamental approach for constructing the TT format is the well-known TT-SVD method, which performs singular value decompositions (SVDs) on the successive matrices sequentially. But in practical applications, it is often unnecessary to compute full SVDs. In this article, we propose a new method called the TT-UTV. It utilizes the virtues of rank-revealing UTV decomposition to compute the TT format for a large-scale tensor, resulting in lower computational cost. We analyze the error bounds on the accuracy of these algorithms in both the URV and ULV cases and then recommend different sweep patterns for these two cases. Based on the theoretical analysis, we also formulate the rank-adaptive algorithms with prescribed accuracy. Numerical experiments on various applications, including magnetic resonance imaging data completion, are performed to illustrate their good performance in practice.
keywords:
@@
Tensor-train decomposition, UTV decomposition , Rank-adaptive algorithms , Data completion.
journal: Nuclear Physics B
[…]
[author1]organization=School of Mathematical Sciences,addressline=Chengdu University of Technology, city=Chengdu, postcode=610059, state=Sichuan, country=China
[…]
[author2]organization=School of Mathematics and Statistics,addressline=Guizhou University, city=Guiyang, postcode=550025, state=Guizhou, country=China
[…]
[author3]organization=School of Mathematical Sciences and Key Laboratory of Mathematics for Nonlinear Sciences,addressline=Fudan University, city=Shanghai, postcode=200433, state=Shanghai, country=China
1 Introduction
@@
Real-world high-dimensional data are usually represented as tensors, such as genetic sequences, video data, hypergraphs, hyperspectral images, and social network interactions [lu2011survey, hore2016tensor, cui2024discrete, velasco2013classification, Koniusz2021, ding2018tensor]. Tensors provide a natural way to model multi-dimensional data and multi-modal nonlinear interactions. However, directly processing these higher-order tensors demands a significant amount of resources, as the time and space complexity grow exponentially with tensor order. This phenomenon is known as the “curse of dimensionality” [Indyk1998]. Fortunately, tensors in the real world often exhibit a low-rank structure, and the tensor-train decomposition [oseledets2011tensor, oseledets2010tt], also known as the matrix product state representation [Perez-Garcia2007MPS], can exploit the inherent low-rank structure to mitigate the curse of dimensionality in many practical scenarios.
The tensor-train (TT) decomposition has found widespread applications in areas such as machine learning [yang2017tensor, novikov2015tensorizing, Sidiropoulos2017], quantum many-body physics [Schollwock2011wave], signal processing [Sedighin2020matrix, miron2020tensor], data completion [xu2023tensor, Steinlechner2016, Cai2022provable], and numerical partial differential equations [Dolgov2021, Richter2021Solving]. For instance, the TT decomposition has been employed to compress deep neural networks while maintaining competitive performance [yang2017tensor, novikov2015tensorizing]. In quantum physics, the TT format enables efficient representation and manipulation of wave functions and operators, thereby significantly reducing computational demands [Schollwock2011wave]. Over the past decade, the TT model has also been adopted for certain high-dimensional data completion problems [Steinlechner2016, Cai2022provable]. In addition, TT-based methods are increasingly applied in solving high-dimensional partial differential equations, where traditional methods often become computationally prohibitive [Dolgov2021, Richter2021Solving], to name a few.
The research on efficient algorithms for constructing TT decomposition of large-scale tensors remains an active area. For a $d$th-order tensor, the TT-SVD algorithm [oseledets2011tensor] sequentially utilizes d1d-1 singular value decompositions (SVDs) to extract core tensors one by one, which serves as a foundational approach to decomposing a higher-order tensor into the TT format. TT-cross approximation [oseledets2010tt] replaces the full SVD computation in the TT-SVD algorithm with skeleton decomposition, making it particularly suitable for large-scale tensors with missing or incomplete data. The alternating least squares method, as an iterative method, computes the TT decomposition by solving a sequence of least squares problems to iteratively refine the TT cores. This method is relatively dependent on the selection of initialization and may exhibit slower convergence for tensors with large condition numbers[Holtz2012the]. In recent years, experts have also proposed parallel algorithms to reduce the runtime for computing tensor-train decomposition; their subroutines still rely on the matrix SVD [zniyed2020tt, shi2023parallel, xie2023error].
@@
In many applications, the SVD is computationally demanding and difficult to update. These limitations have motivated the development of the rank-revealing two-sided orthogonal decomposition, also called the UTV decomposition [Stewart1992an]. The UTV decomposition provides approximately the same numerical rank and subspace information as SVD but at a lower computational cost [Stewart1992an, Fierro1997low, Fierro1995bounding, Demmel2007fast]. Furthermore, UTV decomposition can be efficiently implemented on modern hardware architectures by leveraging block matrix operations that are amenable to parallel processing [Martinsson2019rand]. Due to these virtues, it has become a common alternative to SVD in many applications, including signal processing and machine learning [yang1995projection, Zhou2012stable]. Recently, the UTV decomposition has also been extended to the decompositions of higher-order tensors. For example, a truncated multilinear UTV decomposition [Vandecappelle2022from] is introduced for tensors from the truncated multilinear SVD [DeLathauwer2000a, Vannieuwenhoven2012a]; the t-UTV decomposition [che2022efficient] is proposed for third-order tensors as a substitute for the t-SVD [kilmer2011factorization] based on the tensor t-product.
@@
In this article, we investigate whether the UTV decomposition can be leveraged to construct the TT decomposition as an economic alternative to the TT-SVD. The goal is to save computational costs by utilizing the advantages of UTV decomposition over SVD for large-scale unfolding matrices. This new approach, termed TT-UTV, is designed to operate under both fixed TT-ranks and fixed precision. There exist two types of TT-UTV algorithms corresponding to the URV and ULV cases, respectively. We provide an in-depth error analysis for these algorithms and demonstrate their effectiveness through various numerical experiments, including video recovery and magnetic resonance imaging data completion.
@@
The rest of this paper is organized as follows. Section 2 reviews the necessary preliminary notions. In Section 3, we introduce the TT-ULV algorithm in a left-to-right sweep and TT-URV algorithm in a right-to-left sweep, respectively. We derive rigorous error bounds for them, which explain how the UTV decompositions are used to achieve accurate and efficient TT decomposition, and their computational cost is analyzed. In Section 4, we conduct performance evaluations of the TT-UTV algorithm with TT-SVD in various applications. The conclusions are drawn in Section 5.
2 Preliminaries
@@
Notations. A $d$th-order tensor is a multidimensional array with $d$ free indices. Matrices are second-order tensors, and vectors are first-order tensors. We mainly inherit the usage of notations from the comprehensive works of Kolda and Bader [kolda2009tensor]. The calligraphic capital letters such as $\mathcal{A}$, boldface capital letters such as $\mathbf{A}$, and boldface lowercase letters such as $\mathbf{a}$ are used to denote the tensors, matrices, and vectors, respectively. $\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$ denotes the set of all $d$th-order tensors of dimensions $(I_{1},\dots,I_{d})$ over the real number field $\mathbb{R}$, where $I_{k}$ is the dimension of the $k$th mode of the tensor. The entries of $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$ are accessed by $\mathcal{A}_{i_{1}i_{2}\cdots i_{d}}$ or $\mathcal{A}(i_{1},i_{2},\dots,i_{d})$ for $1\leq i_{k}\leq I_{k}$. For a tensor $\mathcal{A}$, we denote by $\mathbf{A}_{k}$ its $k$th unfolding matrix.
2.1 Tensor basics
The tensor-train (TT) decomposition represents a higher-order tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$ as the contractions of small factor tensors
where $r_{0}=r_{d}=1$, these third-order tensors $\mathcal{G}^{(k)}\in\mathbb{C}^{r_{k-1}\times I_{k}\times r_{k}}$ for $k=1,2,\ldots,d$ are called TT-cores, and $(r_{1},r_{2},\ldots,r_{d-1})$ represent TT-ranks. Fig. 1 is a graphical representation of the TT decomposition of a $d$th-order tensor, with core tensors depicted as blue blocks and indices represented as connecting lines. It looks like a train, as its name implies. In virtue of the TT decomposition, the tensor $\mathcal{A}\in\mathbb{C}^{I_{1}\times I_{2}\times\cdots\times I_{k}}$ can be compressed into ${\textstyle\sum_{k=1}^{d}}r_{k-1}I_{k}r_{k}$ parameters, which becomes linearly dependent on the tensor order $d$. Lower TT-ranks imply lower memory consumption and computational costs.
Fig. 1: Graphical illustration of the TT decomposition for a $d$th-order tensor.
@@ 2.1 Tensor basics
Tensor unfolding, the process of rearranging the elements of a tensor into a matrix or vector, is a frequently used operation in tensor analysis and numerical computations [kolda2009tensor]. The reverse lexicographic ordering (as the ‘reshape’ command in MATLAB) is one of the most commonly used approaches for integrating multiple indices into a single long index. Concretely, for the tensor dimensions $\mathbf{I}=\left(I_{1},\dots,I_{d}\right)$ and a corresponding index vector $\mathbf{i}=\left(i_{1},\dots,i_{d}\right)$, the reverse lexicographic order map
$\mathrm{ivec}(\mathbf{i},\mathbf{I}):=i_{1}+\sum_{k=2}^{d}(i_{k}-1)\prod_{j=1}^{k-1}I_{j}.$ (2)
Definition 2.1 (The $k$th unfolding matrix [oseledets2011tensor]).
@@
For a $d$th-order tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$, the matrix $\mathbf{A}_{k}\in\mathbb{R}^{(I_{1}\cdots I_{k})\times(I_{k+1}\cdots I_{d})}$ is called the $k$th unfolding matrix of $\mathcal{A}$ such that
@@
where $i=\mathrm{ivec}\left((i_{1},\dots,i_{k}),(I_{1},\dots,I_{k})\right)$ and $j=\mathrm{ivec}\left((i_{k+1},\dots,i_{d}),(I_{k+1},\dots,I_{d})\right)$. That is to say, the first $k$ indices and the last dkd-k indices of $\mathcal{A}$ are fused into the row and column indices of the unfolding matrix $\mathbf{A}_{k}$, respectively, following the reverse lexicographic order $\rm ivec$ (2).
Definition 2.2 (Orthogonal TT-cores [oseledets2011tensor]).
The TT-cores $\{\mathcal{G}^{(k)},k=1,\dots,d\}$ in Eq. (1) are called left-orthogonal if for $k=1,\dots,d-1$,
@@
where $\mathbf{G}^{(k)}_{2}\in\mathbb{R}^{r_{k-1}I_{k}\times r_{k}}$ are the second unfolding matrices of the TT-cores $\mathcal{G}^{(k)}$. They are right-orthogonal if for $k=2,\dots,d$,
$\mathbf{G}^{(k)}_{1}\mathbf{G}^{(k)\top}_{1}=\mathbf{I}_{r_{k-1}},$
where $\mathbf{G}^{(k)}_{1}\in\mathbb{R}^{r_{k-1}\times I_{k}r_{k}}$ are the first unfolding matrices of the TT-cores $\mathcal{G}^{(k)}$, $\mathbf{I}_{r_{k}}\in\mathbb{R}^{r_{k}\times r_{k}}$ denotes the identity matrix.
The TT-SVD algorithm computes the TT format (1) with left-orthogonal cores by a single left-to-right sweep or the right-orthogonal cores by a right-to-left sweep; more details can be found in [oseledets2011tensor].
Definition 2.3 (Mode product [kolda2009tensor]).
The mode-$k$ product of $\mathcal{X}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$ with a matrix $\mathbf{A}\in\mathbb{R}^{J\times I_{k}}$ is an $I_{1}\times\cdots\times I_{k-1}\times J\times I_{k+1}\times\cdots\times I_{d}$ tensor, given by
@@
Apart from the tensor-train decomposition, the Tucker format [tucker1966some] is another compact form for large-scale tensors, which represents a $d$th-order tensor via the mode products of a smaller-scaled core tensor and $d$ factor matrices:
@@
For $\mathbf{A}\in\mathbb{R}^{m\times n}$ and $\mathbf{B}\in\mathbb{R}^{p\times q}$, the Kronecker product $\mathbf{A}\otimes\mathbf{B}$ is the $mp\times nq$ block matrix
@@ 2.2 UTV decompositions
$\mathbf{A}=\mathbf{U}\begin{bmatrix}\mathbf{L}_{11}&\mathbf{O}_{r\times(n-r)}\\ \mathbf{L}_{21}&\mathbf{L}_{22}\end{bmatrix}\mathbf{V}^{\top}.$ (4)
Here $\mathbf{L}_{11}$ and $\mathbf{L}_{22}$ are lower triangular, $\mathbf{L}_{21}\in\mathbb{R}^{(m-r)\times r}$ with small entries. Moreover, suppose the numerical rank of $\mathbf{A}$ is $r$, then the ULV decomposition is said to be rank-revealing if
@@
where $\mathbf{R}_{11}\in\mathbb{R}^{r\times r}$, $\mathbf{R}_{22}\in\mathbb{R}^{(m-r)\times(n-r)}$ are upper triangular, $\mathbf{O}$ denotes the zero matrix whose elements are all 0, and $\mathbf{R}_{12}\in\mathbb{R}^{r\times(n-r)}$ contains small entries. The URV decomposition is said to be rank-revealing if
$\sigma_{\rm min}(\mathbf{R}_{11})=O(\sigma_{r})\quad\mathrm{and}\quad\left\|\begin{bmatrix}\mathbf{R}_{12}^{\top}&\mathbf{R}_{22}^{\top}\end{bmatrix}\right\|_{2}=O(\sigma_{r+1}).$ (7)
The smaller the norm of the off-diagonal block, the better the truncated low-rank approximations in Eqs. (4) and (6). In fact, the UTV decomposition of a matrix is not unique, and various matrix decompositions proposed in the literature can be regarded as UTV-type, such as the rank-revealing QR decomposition [CHAN198767], the pivoted QLP decomposition [stewart1999qlp], and the SVD, which is also a special case of UTV decomposition when the middle matrix is diagonal. There are many efficient algorithms for UTV decomposition. Stewart [Stewart1992an] introduced the algorithms for UTV decomposition that are well-suited for high-rank matrices. Fierro etal.et\ al. [fierro1999utv, Fierro1997low] proposed the low-rank UTV algorithms and built a Matlab package named ‘UTV Tools’ for efficient rank-revealing decompositions. In recent years, more fast randomized algorithms have been proposed for UTV decomposition, which are flexible for parallel processing in modern architectures [Martinsson2019rand, Lee2018rankrev, kaloorazi2020efficient]. In many practical applications, they can achieve comparable accuracy at a lower computational cost than classical SVD.
3 The proposed methods
In this section, we propose the TT-UTV algorithms for efficiently computing the approximate TT format by sequentially decomposing the cores via UTV factorizations. We discuss two variants of the algorithm, corresponding to the ULV and URV decompositions, respectively. The feasibility of the proposed methods is rigorously supported by theoretical analysis, based on which we recommend distinct update strategies for each case and further introduce fixed-precision algorithms that can adaptively determine the TT-ranks.
3.1 TT-ULV case